OpenStudy (anonymous):

Suppose \[f : R^{n} \rightarrow R\]is differentiable and ||f(t)|| = 1 for all t. Show that the dot product of Df(t) and f(t) is always 0. Describe what this means geometrically.

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

any ideas?

ganeshie8 (ganeshie8):

Lets think about geometry first

ganeshie8 (ganeshie8):

when n=2, \(\|f(t)\|=1\) gives us circular motion : 

OpenStudy (anonymous):

so in Rn it would be a ball?

ganeshie8 (ganeshie8):

yeah we can say the motion is \(n\) dimensional ball

OpenStudy (anonymous):

then how can we calculate the Df and f dot product?

OpenStudy (anonymous):

any leads?

OpenStudy (perl):

see if you can show it is true for n = 3, then try to generalize it

OpenStudy (perl):



OpenStudy (perl):

woops, use f(t) instead of r(t)

OpenStudy (anonymous):

how do i put these into english words tho. I suck at english. lol

OpenStudy (perl):

$$ \Large || A ||^2 =A ~. A\\ \\ \therefore \\ \Large ||f(t)||^2 = f(t)~.f(t) $$

OpenStudy (anonymous):

what about Df(t) ? it's Rn - R so it's gonna be a 1xN matrix

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

still not getting it

OpenStudy (perl):

the left side is zero, because its a derivative of a constant $$ $$

OpenStudy (anonymous):

since f(t) and f'(t) the the line and it's perpendicular line, shouldn't be always have a dot product of zero? Since A dot B = |A||B| cos , and the cos is always gonna be pi/2

OpenStudy (perl):

$$ \Large || \vec{a} ||^2 =\vec{a} ~. \vec{a}\\ \\ \therefore \\ \Large ||f(t)||^2 = f(t)~.f(t)\\ \Large ||f(t)|| = 1 \\ \\ \therefore\\ \Large 0 =\frac{d}{dt}(1^2) = \frac{d}{dt}(||f(t)||^2) = (f(t)~.f(t)) ~' \\ \Large = { f'(t) . f(t) + f(t) . f ' (t)\\ = 2 f ' (t) . f(t) \\ \therefore \\ \Large 0 = 2 f ' (t) . f(t) \\ \therefore \\ \Large 0 = f ' (t) . f(t) } $$

OpenStudy (perl):

yes but how do you know that f(t) and f ' (t) are perpendicular

OpenStudy (perl):

" since f(t) and f'(t) `the the line` and it's perpendicular line, shouldn't be always have a dot product of zero? Since A dot B = |A||B| cos , and the cos is always gonna be pi/2" what did you mean here?

OpenStudy (anonymous):

oh ok. I might confused it with some sort of lemma that can be used directly.

OpenStudy (anonymous):

I meant " is the line and ......"

OpenStudy (perl):

I drew r ' (t) perpindicular to r(t) in the picture, but that is what we want to prove. it is true that r ' (t) is tangent to the circle , but its not obvious that r ' (t) is perpindicular to r (t) .

OpenStudy (anonymous):

got you. Thank you ! You are my favourite

OpenStudy (perl):

This result is consistent with elementary geometry, but it is better not to use elementary geometry. Elementary geometry will only tell us about at most 2 or 3 dimensions, one dimension it is trivially true. Vector algebra goes to n dimensions , so it is more general and more powerful. Now there might be a way to prove this with the cos theta definition of dot product, but I don't see a simple proof.

OpenStudy (anonymous):

yea, anyway, thank you for all the help and your time.XD

OpenStudy (perl):

Note that this is a vector valued function, so we used a shorthand: $$ \Large f(t)=(f_1(t),f_2(t),\ldots,f_n(t)) $$

OpenStudy (perl):

otherwise it looks like a 1 valued function